DENSITY AND TEMPERATURE STUDY OF THE RAMAN NON-COINCIDENCE EFFECT IN ALKYL BENZOATES

\underline{V. L. SLAGER}, H.-C. CHANG, and J. JONAS, Department of Chemistry, School of Chemical Sciences, and Beckman Institute for Advanced Science and Technology, University of Illinois, Urbana, IL 61801.

The non-coincidence (Δν) of the isotropic and anisotropic peak frequencies of the Raman carbonyl band was studied for a series of alkyl benzoates. Methyl benzoate, ethyl benzoate, propyl benzoate, butyl benzoate, hexyl benzoate and 2-ethylhexyl benzoate were observed over a density range of 0.95 to 1.20 g cm^{−3} at 20 and 40 ^{∘}C. The characteristic density dependence of the value of the non-coincidence effect for each alkyl benozate correlates well with the length and branching of the alkyl side chain. In fact, it was possible to interpret the density dependence of Δν in terms of packing fraction and conformational changes of the alkyl side chain.